# MoC-System: Efficient Fault Tolerance for Sparse Mixture-of-Experts Model Training

### v2

@@ 1. Introduction
Transformer-based large language models (LLMs), which scale to billions or even trillions of parameters, have emerged as the most trending topic in AI research due to their impressive capabilities (achiam2023gpt, ; chowdhery2023palm, ; wei2022chain, ; ouyang2022training, ; wei2022emergent, ; brown2020language, ; attention, ). Recently, the sparsely-gated mixture-of-experts (MoE) has become the preferred methodology to increase parameter counts and enhance the model quality of LLMs without a proportional increase in computational requirements (ShazeerMMDLHD17, ; ViTMoE, ; riquelme2021scaling, ; jiang2024mixtral, ; cai2024survey, ). To facilitate the training of MoE models and their deployment across expansive computing clusters, distributed training systems have been refined to incorporate expert parallelism (EP) (lepikhin2020gshard, ; fedus2022switch, ; hwang2023tutel, ; singh2023hybrid, ; he2021fastmoe, ; cai2024shortcut, ) alongside established frameworks of data parallelism (DP) (rajbhandari2020zero, ; ren2021zero, ) and model parallelism (shoeybi2019megatron, ; narayanan2021efficient, ; smith2022using, ; huang2019gpipe, ; narayanan2019pipedream, ; li2021sequence, ; korthikanti2023reducing, ). With the escalation in the number of deployed computing devices and the incidence of faults (maeng2021understanding, ; wu2023transom, ; jiang2024megascale, ; gupta2017failures, ; di2014lessons, ), ensuring fault tolerance has become a critical component of AI system infrastructure.
@@
Although numerous studies have effectively addressed fault tolerance for dense (non-MoE) models through periodical checkpoints (wang2023gemini, ; jiang2024megascale, ; nicolae2020deepfreeze, ; wu2023transom, ; maeng2021understanding, ; mohan2021checkfreq, ), the distinctive characteristics of MoE models necessitate specialized strategies to assure their reliable and efficient fault-tolerant training. As MoE models scale to unprecedented sizes, the primary challenge in fault tolerance is the substantial increase in checkpoint size, which poses a storage burden that distributed filesystems struggle to handle efficiently (rajbhandari2022deepspeed, ; shen2022se, ; du2022glam, ). Even with prevailing sharded and asynchronous checkpointing strategies (nicolae2020deepfreeze, ; mohan2021checkfreq, ; wang2024fastpersist, ; maurya2024datastates, ; chen2023cost, ), the enlarged checkpointing duration cannot be fully overlapped with the training process, yielding additional costs to the total training time.
@@
Pioneering efficient fault tolerance for MoE model training, we introduce the Mixture-of-Checkpoint System (MoC-System). The name ”Mixture-of-Checkpoint” reflects the system’s design to orchestrate the vast array of checkpoint shards produced during distributed training. MoC-System features an innovative Partial Experts Checkpointing (PEC) mechanism, an algorithm-system co-design that reduces the checkpoint size for MoE models by selectively saving only a subset of experts. Specifically, PEC selectively saves $K_{pec}$ experts per MoE layer during each checkpointing, while fully saving the non-expert parameters of the model.
@@
It is inspired by observations in MoE model fine-tuning, where updating only the non-expert parameters can achieve the same accuracy as updating all parameters while updating only the expert parameters leads to a drastic reduction in model accuracy (zoph2022st, ). This is believed to be due to the sparsity of the MoE structure, which makes it less sensitive to a limited number of updates, as supported by observations that MoE models generally require larger volumes of pre-training data (artetxe2021efficient, ; fedus2022switch, ; xue2022go, ). Building on existing algorithm-system co-design efforts that leverage LLMs’ features to optimize computation (dao2022flashattention, ; liu2024efficient, ), communication (rajbhandari2022deepspeed, ; he2022fastermoe, ; wang2024towards, ), and memory (yi2023edgemoe, ; hwang2024pre, ), we innovatively apply the co-design method to fault tolerance.
@@
We implement the MoC-System and conduct experiments upon the Megatron-DeepSpeed (Megatron_DeepSpeed, ; smith2022using, ; rajbhandari2022deepspeed, ), which is an acclaimed open-source framework supporting the predominant MoE training strategy of ZeRO-2 DP (rajbhandari2020zero, ) and EP. Our experimental results from training the GPT-350M-16E model demonstrate that the PEC approach achieves a 57.7% reduction in the total checkpoint size. Furthermore, recovery from PEC checkpoints maintains comparable validation loss during pre-training and even achieves an average accuracy increase of 1.08% on downstream tasks. Additionally, with all optimizations applied, MoC-System reduces overhead for each checkpointing process by up to 98.9% compared to the original method, and speeds up each training iteration with checkpointing by up to 5.12$\times$.
@@ 2.1. Sparse Mixture-of-Experts (MoE) Models
The sparse Mixture-of-Experts (MoE) layer (ShazeerMMDLHD17, ), consists of multiple feed-forward networks (FFNs), termed “experts”, and a trainable gating network for selectively activating a subset of these experts. Formally, with $N$ expert networks $\{E_{i}\}_{1}^{N}$, gating network $G$, and input $x$, the MoE layer’s output can be formulated as:
@@ 2.2. Distributed Training of MoE Models
The adoption of MoE in LLMs introduces new challenges to existing training and inference systems, due to its inherently sparse and dynamic computational workload. GShard (lepikhin2020gshard, ) pioneers the parallel strategy of Expert Parallelism (EP) by facilitating parallel gating and expert computation. Specifically, EP assigns distinct experts to each distributed computing device such as GPU and TPU, and passes input tokens to the corresponding experts via All-to-All communication. Following this, EP has ascended as a pivotal strategy, enabling the efficient scaling of MoE model training (fedus2022switch, ; rajbhandari2022deepspeed, ; hwang2023tutel, ; singh2023hybrid, ).
As depicted in Figure 1 (a), EP can be viewed as an augmentation of Data Parallelism (DP) (rajbhandari2020zero, ; ren2021zero, ; rajbhandari2021zero, ), where each expert within an MoE layer is allocated to a distinct DP rank (e.g., “Expert1-0” on “Rank0” and “Expert1-1” on “Rank1”), while all non-expert layers (e.g., “Atten0”, “FFN0”, and “Atten1”) are replicated across DP ranks. Moreover, the synergy of EP with other parallel strategies, such as Tensor Parallelism (TP) (shoeybi2019megatron, ; smith2022using, ; narayanan2021efficient, ), Pipeline Parallelism (PP) (huang2019gpipe, ; narayanan2019pipedream, ; qi2023zero, ), has been explored to enhance the scalability and efficiency of MoE model training in expansive distributed settings (fedus2022switch, ; rajbhandari2022deepspeed, ; singh2023hybrid, ; hwang2023tutel, ; zhai2023smartmoe, ; he2022fastermoe, ; wei2024skywork, ). From the viewpoint of checkpointing, a notable distinction between EP and other parallel strategies is EP’s flexibility in distributing diverse parameters across DP ranks. In contrast, TP and PP maintain parameters replicated across all DP ranks, limiting their adaptability within each DP rank.
In this work, we primarily focus on distributed training with the hybrid parallel strategy of ZeRO-2 DP + EP (notably, ZeRO-1 is analogous to ZeRO-2 from the view of checkpointing (rajbhandari2020zero, )), which has emerged as the predominant approach for training MoE models (rajbhandari2022deepspeed, ; llama-moe-2023, ; wei2024skywork, ). This approach is highlighted for its accessibility and efficiency, supported by Megatron-DeepSpeed (Megatron_DeepSpeed, ; smith2022using, ), an acclaimed open-source distributed training framework. Moreover, extensive practical experience with large-scale distributed systems has demonstrated its superior performance, minimizing communication overhead while remaining memory-efficient (rajbhandari2020zero, ; rajbhandari2022deepspeed, ; cai2024survey, ; SimAI, ). Additionally, our proposed checkpointing techniques can be seamlessly extended to other hybrid parallel strategies, encompassing TP and PP, as they can be viewed as the modularity of each DP rank.
@@ 2.3. Fault-tolerant Checkpointing for Distributed Training System
Checkpoint serves as a critical mechanism for augmenting fault tolerance in distributed training systems by facilitating the periodic preservation and recovery of model states during training (mohan2021checkfreq, ; jiang2024megascale, ; wu2023transom, ; le2023bloom, ; koo1987checkpointing, ). As illustrated in Figure 2, the saved model states at each checkpoint comprise learnable model parameters for the expert part (12% of the total volume) and that for the non-expert part (2%), optimizer states for the expert part (74%) and that for the non-expert part (12%), along with other crucial states (less than 1%), such as epoch/iteration numbers and Random Number Generator states. The checkpoint ensures that training progress is not lost in unexpected faults and can be recovered after a restart.
However, the checkpointing process incurs significant data transfer and storage overhead, alongside additional overhead in the event of a fault. The total overhead introduced by fault tolerance with checkpoint during the entire model training, denoted as $O_{ckpt}$, can be quantified by aggregating the overhead of model states saving $O_{save}$ during normal training, the overhead of system/task restart $O_{restart}$ and lost training progress $O_{lost}$ when a fault occurs, as illustrated in Figure 2. It is formulated as:
@@
As demonstrated in Figure 3, checkpointing model states from GPU memory to distributed persistent storage involves two phases: transferring from GPU memory to CPU memory (GPU-to-CPU snapshot) and from CPU memory to distributed persistent storage (CPU-to-Storage persist). The CPU-to-Storage persist phase involves serializing the model states and writing them to a distributed filesystem via the network, while the GPU-to-CPU snapshot phase copies tensors through PCIe. Given that both phases can significantly hinder training progress if executed in a blocking manner, asynchronously processing and overlapping them with ongoing training has emerged as a critical method to enhance checkpointing efficiency (mohan2021checkfreq, ; wang2023reliable, ; jiang2024megascale, ; wan2024bytecheckpoint, ; wang2024fastpersist, ; maurya2024datastates, ; chen2023cost, ).
@@
Considering that data volume determines the duration of communication and storage, eliminating redundancies and reducing checkpoint size through data-parallel sharding (nicolae2020deepfreeze, ; wang2023reliable, ; wan2024bytecheckpoint, ; lian2024universal, ) is an effective optimization. Since the original DP replicates model states across all the DP ranks, each rank can store a distinct sharding of the states, collectively forming the complete model states through the aggregation of all ranks’ shards, as depicted in Figure 3.
With the evolution of DP techniques, model states may already be uniformly partitioned across each DP rank. For instance, ZeRO-1 and ZeRO-2 DP (rajbhandari2020zero, ) partition the optimizer states, whereas ZeRO-3 DP and Fully Sharded Data Parallel (FSDP) (zhao2023pytorch, ) partition both the model parameters and optimizer states. As discussed in Section 2.2, our focus is on the ZeRO-2 DP + EP scenario, where model parameters are replicated across each DP rank, as illustrated in Figure 1.
However, existing distributed training frameworks lack an efficient data-parallel sharding strategy for MoE model training. For instance, the Megatron-DeepSpeed framework (Megatron_DeepSpeed, ) confines the checkpointing of expert model states to the first EP group, as illustrated by Figure 7 (a), neglecting the potential of distributed sharding across all EP groups. In contrast, we implement fully sharded checkpointing for MoE model training and further introduce an adaptive sharding strategy with our PEC mechanism, outperforming the commonly used equal sharding strategy.
@@
Due to the superior bandwidth of GPU-to-CPU copy and compute network data transfers compared to CPU-to-Storage persist, several studies minimize $O_{save}$ by opting to store model states in the CPU memory of other nodes instead of persistent storage (wang2023gemini, ; wang2023reliable, ). This approach significantly reduces the duration of checkpointing, thereby allowing for lower $I_{ckpt}$ and $O_{ckpt}$.
However, the in-memory checkpointing solution encounters reliability problems within real-world large-scale GPU clusters (wan2024bytecheckpoint, ). In such environments, multiple nodes within the same backup group may experience simultaneous failures, resulting in severe data loss. In contrast, we introduce a two-level checkpoint management strategy that benefits from the efficiency of in-memory checkpointing while ensuring fault tolerance across a wide range of scenarios.
@@
Previous research (qiao2019fault, ) has demonstrated that the iterative-convergent nature of machine learning (ML) training is capable of compensating for the inconsistencies introduced by partial checkpointing and recovery to some extent on Parameter Server (PS) distributed training scenarios. Given that the model parameters are distributed across multiple PS nodes, a partial failure of these nodes is likely to result in only a partial loss of the updates to the model parameters. Compared to the process of saving and reloading the entire model states, partial checkpointing and recovery strategies can significantly decrease the data volume required for checkpoints. This approach has subsequently proven to be effective in the training of Deep Learning Recommendation Model (DLRM) (naumov2019deeplearningrecommendationmodel, ; maeng2021understanding, ; eisenman2022check, ), which only access and update a small segment of the model in each iteration.

### v3

@@ Abstract.
Fault Tolerance, Checkpoint, Mixture of Experts, Large Language Models, Training
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 30–April 3, 2025; Rotterdam, Netherlands.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsisbn: 979-8-4007-1079-7/25/03doi: 10.1145/3676641.3716006ccs: Computer systems organization Reliability
@@ 1. Introduction
Transformer-based large language models (LLMs), which scale to billions or even trillions of parameters, have emerged as the most trending topic in AI research due to their impressive capabilities (Achiam et al., 2023; Chowdhery et al., 2023; Wei et al., 2022b; Ouyang et al., 2022; Wei et al., 2022a; Brown et al., 2020; Vaswani et al., 2017). Recently, the sparsely-gated mixture-of-experts (MoE) has become the preferred methodology to increase parameter counts and enhance the model quality of LLMs without a proportional increase in computational requirements (Shazeer et al., 2016; Riquelme et al., 2021; Jiang et al., 2024b; Cai et al., 2024b). To facilitate the training of MoE models and their deployment across expansive computing clusters, distributed training systems have been refined to incorporate expert parallelism (EP) (Lepikhin et al., 2020; Fedus et al., 2022; Hwang et al., 2023; Singh et al., 2023; He et al., 2021; Cai et al., 2024a) alongside established frameworks of data parallelism (DP) (Rajbhandari et al., 2020; Ren et al., 2021) and model parallelism (Shoeybi et al., 2019; Narayanan et al., 2021; Smith et al., 2022; Huang et al., 2019; Narayanan et al., 2019; Li et al., 2021; Korthikanti et al., 2023). With the escalation in the number of deployed computing devices and the incidence of faults (Maeng et al., 2021; Wu et al., 2023; Jiang et al., 2024a; Gupta et al., 2017; Di Martino et al., 2014), ensuring fault tolerance has become a critical component of AI system infrastructure.
@@
Although numerous studies have effectively addressed fault tolerance for dense (non-MoE) models through periodical checkpoints (Wang et al., 2023a; Jiang et al., 2024a; Nicolae et al., 2020; Wu et al., 2023; Maeng et al., 2021; Mohan et al., 2021), the distinctive characteristics of MoE models necessitate specialized strategies to assure their reliable and efficient fault-tolerant training. As MoE models scale to unprecedented sizes, the primary challenge in fault tolerance is the substantial increase in checkpoint size, which poses a storage burden that distributed filesystems struggle to handle efficiently (Rajbhandari et al., 2022; Shen et al., 2022; Du et al., 2022). Even with prevailing sharded and asynchronous checkpointing strategies (Nicolae et al., 2020; Mohan et al., 2021; Wang et al., 2024a; Maurya et al., 2024; Chen et al., 2023), the enlarged checkpointing duration cannot be fully overlapped with the training process, yielding additional costs to the total training time.
@@
Pioneering efficient fault tolerance for MoE model training, we introduce the Mixture-of-Checkpoint System (MoC-System). The name “Mixture-of-Checkpoint” reflects the system’s design to orchestrate the vast array of checkpoint shards produced during distributed training. MoC-System features an innovative Partial Experts Checkpointing (PEC) mechanism, an algorithm-system co-design that reduces the checkpoint size for MoE models by selectively saving only a subset of experts. Specifically, PEC selectively saves $K_{pec}$ experts per MoE layer during each checkpointing, while fully saving the non-expert parameters of the model.
@@
It is inspired by observations in MoE model fine-tuning, where updating only the non-expert parameters can achieve the same accuracy as updating all parameters while updating only the expert parameters leads to a drastic reduction in model accuracy (Zoph et al., 2022). This is believed to be due to the sparsity of the MoE structure, which makes it less sensitive to a limited number of updates, as supported by observations that MoE models generally require larger volumes of pre-training data (Artetxe et al., 2021; Fedus et al., 2022; Xue et al., 2022). Building on existing algorithm-system co-design efforts that leverage LLMs’ features to optimize computation (Dao et al., 2022; Liu et al., 2024b), communication (Rajbhandari et al., 2022; He et al., 2022; Wang et al., 2024b), and memory (Yi et al., 2023; Hwang et al., 2024), we innovatively apply the co-design method to fault tolerance.
@@
We implement the MoC-System and conduct experiments upon the Megatron-DeepSpeed (Microsoft, 2022; Smith et al., 2022; Rajbhandari et al., 2022), which is an acclaimed open-source framework supporting the predominant MoE training strategy of ZeRO-2 DP (Rajbhandari et al., 2020) and EP. Our experimental results from training the GPT-350M-16E model demonstrate that the PEC approach achieves a 57.7% reduction in the total checkpoint size. Furthermore, recovery from PEC checkpoints maintains comparable validation loss during pre-training and even achieves an average accuracy increase of 1.08% on downstream tasks. Additionally, with all optimizations applied, MoC-System reduces overhead for each checkpointing process by up to 98.9% compared to the original method, and speeds up each training iteration with checkpointing by up to 5.12$\times$.
@@ 2.1. Sparse Mixture-of-Experts (MoE) Models
The sparse Mixture-of-Experts (MoE) layer (Shazeer et al., 2016), consists of multiple feed-forward networks (FFNs), termed “experts”, and a trainable gating network for selectively activating a subset of these experts. Formally, with $N$ expert networks $\{E_{i}\}_{1}^{N}$, gating network $G$, and input $x$, the MoE layer’s output can be formulated as:
@@ 2.2. Distributed Training of MoE Models
The adoption of MoE in LLMs introduces new challenges to existing training and inference systems, due to its inherently sparse and dynamic computational workload. GShard (Lepikhin et al., 2020) pioneers the parallel strategy of Expert Parallelism (EP) by facilitating parallel gating and expert computation. Specifically, EP assigns distinct experts to each distributed computing device such as GPU and TPU, and passes input tokens to the corresponding experts via All-to-All communication. Following this, EP has ascended as a pivotal strategy, enabling the efficient scaling of MoE model training (Fedus et al., 2022; Rajbhandari et al., 2022; Hwang et al., 2023; Singh et al., 2023).
As depicted in Figure LABEL:fig:DP_EP_overview_a, EP can be viewed as an augmentation of Data Parallelism (DP) (Rajbhandari et al., 2020; Ren et al., 2021; Rajbhandari et al., 2021), where each expert within an MoE layer is allocated to a distinct DP rank (e.g., “Expert1-0” on “Rank0” and “Expert1-1” on “Rank1”), while all non-expert layers (e.g., “Atten0”, “FFN0”, and “Atten1”) are replicated across DP ranks. Moreover, the synergy of EP with other parallel strategies, such as Tensor Parallelism (TP) (Shoeybi et al., 2019; Smith et al., 2022; Narayanan et al., 2021), Pipeline Parallelism (PP) (Huang et al., 2019; Narayanan et al., 2019; Qi et al., 2023), has been explored to enhance the scalability and efficiency of MoE model training in expansive distributed settings (Fedus et al., 2022; Singh et al., 2023; Hwang et al., 2023; Zhai et al., 2023; He et al., 2022; Wei et al., 2024). From the checkpoint perspective, a notable distinction between EP and other parallelism is EP’s flexibility in distributing diverse parameters across DP ranks. In contrast, TP and PP maintain parameters replicated across all DP ranks, limiting their adaptability within each DP rank.
In this work, we primarily focus on distributed training with the hybrid parallel strategy of ZeRO-2 DP + EP (notably, ZeRO-1 is analogous to ZeRO-2 from the view of checkpointing (Rajbhandari et al., 2020)), which has emerged as the predominant approach for training MoE models (Rajbhandari et al., 2022; Team, 2023; Wei et al., 2024). This approach is highlighted for its accessibility and efficiency, supported by Megatron-DeepSpeed (Microsoft, 2022; Smith et al., 2022), an acclaimed open-source distributed training framework. Moreover, extensive practical experience with large-scale distributed systems has demonstrated its superior performance, minimizing communication overhead while remaining memory-efficient (Rajbhandari et al., 2020, 2022; Cai et al., 2024b; Cloud, 2024). Additionally, our proposed checkpointing techniques can be seamlessly extended to other hybrid parallel strategies, encompassing TP and PP, as they can be viewed as the modularity of each DP rank.
@@ 2.3. Fault-tolerant Checkpointing for Distributed Training System
Checkpoint serves as a critical mechanism for augmenting fault tolerance in distributed training systems by facilitating the periodic preservation and recovery of model states during training (Mohan et al., 2021; Jiang et al., 2024a; Wu et al., 2023; Le Scao et al., 2023; Koo and Toueg, 1987). As illustrated in Figure 3, the saved model states at each checkpoint comprise learnable model parameters for the expert part (12% of the total volume) and that for the non-expert part (2%), optimizer states for the expert part (74%) and that for the non-expert part (12%), along with other crucial states (less than 1%), such as epoch/iteration numbers and Random Number Generator states. The checkpoint ensures that training progress is not lost in unexpected faults and can be recovered after a restart.
However, the checkpointing process incurs significant data transfer and storage overhead, alongside additional overhead in the event of a fault. The total overhead introduced by fault tolerance with checkpoint during the entire model training, denoted as $O_{ckpt}$, can be quantified by aggregating the overhead of a checkpointing process (saving model states) $O_{save}$ during normal training, the overhead of system/task restart $O_{restart}$ and lost training progress $O_{lost}$ when a fault occurs, as illustrated in Figure 3. It is formulated as:
@@
As demonstrated in Figure 4, checkpointing model states from GPU memory to distributed persistent storage involves two phases: transferring from GPU memory to CPU memory (GPU-to-CPU snapshot) and from CPU memory to distributed persistent storage (CPU-to-Storage persist). The CPU-to-Storage persist phase involves serializing the model states and writing them to a distributed filesystem via the network, while the GPU-to-CPU snapshot phase copies tensors through PCIe. Given that both phases can significantly hinder training progress if executed in a blocking manner, asynchronously processing and overlapping them with ongoing training has emerged as a critical method to enhance checkpointing efficiency (Mohan et al., 2021; Wang et al., 2023b; Jiang et al., 2024a; Wan et al., 2024; Wang et al., 2024a; Maurya et al., 2024; Chen et al., 2023).
@@
Considering that data volume determines the duration of communication and storage, eliminating redundancies and reducing checkpoint size through data-parallel sharding (Nicolae et al., 2020; Wang et al., 2023b; Wan et al., 2024; Lian et al., 2024) is an effective optimization. Since the original DP replicates model states across all the DP ranks, each rank can store a distinct sharding of the states, collectively forming the complete model states through the aggregation of all ranks’ shards, as depicted in Figure 4.
With the evolution of DP techniques, model states may already be uniformly partitioned across each DP rank. For instance, ZeRO-1 and ZeRO-2 DP (Rajbhandari et al., 2020) partition the optimizer states, whereas ZeRO-3 DP and Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) partition both the model parameters and optimizer states. As discussed in Section 2.2, our focus is on the ZeRO-2 DP + EP scenario, where model parameters are replicated across each DP rank, as illustrated in Figure 2.
However, existing distributed training frameworks lack an efficient data-parallel sharding strategy for MoE model training. For instance, the Megatron-DeepSpeed framework (Microsoft, 2022) confines the checkpointing of expert model states to the first EP group, as illustrated by Figure LABEL:fig:fully_sharded_checkpointing_a, neglecting the potential of distributed sharding across all EP groups. In contrast, we implement fully sharded checkpointing for MoE model training and further introduce an adaptive sharding strategy with our PEC mechanism, outperforming the commonly used equal sharding strategy.
@@
Due to the superior bandwidth of GPU-to-CPU copy and compute network data transfers compared to CPU-to-Storage persist, several studies minimize $O_{save}$ by opting to store model states in the CPU memory of other nodes instead of persistent storage (Wang et al., 2023a, b). This approach significantly reduces the duration of checkpointing, thereby allowing for lower $I_{ckpt}$ and $O_{ckpt}$.
However, the in-memory checkpointing solution encounters reliability problems within real-world large-scale GPU clusters (Wan et al., 2024). In such environments, multiple nodes within the same backup group may experience simultaneous failures, resulting in severe data loss. In contrast, we introduce a two-level checkpoint management strategy that benefits from the efficiency of in-memory checkpointing while ensuring fault tolerance across a wide range of scenarios.
@@
Previous research (Qiao et al., 2019) has demonstrated that the iterative-convergent nature of machine learning (ML) training is capable of compensating for the inconsistencies introduced by partial checkpointing and recovery to some extent on Parameter Server (PS) distributed training scenarios. Given that the model parameters are distributed across multiple PS nodes, a partial failure of these nodes is likely to result in only a partial loss of the updates to the model parameters. Compared to the process of saving and reloading the entire model states, partial checkpointing and recovery strategies can significantly decrease the data volume required for checkpoints. This approach has subsequently proven to be effective in the training of the Deep Learning Recommendation Model (DLRM) (Naumov et al., 2019; Maeng et al., 2021; Eisenman et al., 2022), which only accesses and updates a small segment of the model in each iteration.